OpenStudy (deannamarx01):

simplify (sin theta - cos theta) - (sin theta + cos theta)^2 @bmk614

OpenStudy (bmk614):

Distribute the exponent. You should get. \[[\sin \theta-\cos \theta]-[\sin^2 \theta+2\sin \theta \cos \theta+\cos^2 \theta]\]

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\)\(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)

OpenStudy (bmk614):

Given sin^2+cos^2=1, You can replace that with 1. You should get \[\sin \theta-\cos \theta-(1+2\sin \theta \cos \theta)\]